# Generative Topological Networks

### v2

@@ 4 Related work
GTNs are related in concept to NFs which seek to map a source distribution into a target distribution using a sequence of bijective transformations. These are typically implemented by a neural network, often in the form of at least tens of neural network blocks (Xu and Campbell 2024) and sometimes more (Xu, Chen, and Campbell 2023). The core of NFs is based on each of these transformations being a diffeomorphism – a specific type of homeomorphism that is more constrained as it is defined between smooth manifolds (as opposed to any topological spaces, including any manifolds) and requires that the function and its inverse are differentiable (as opposed to just continuous for general homeomorphisms). NFs also require that the log-determinant of the Jacobian of these transformations is tractable, posing a constraint on the model architecture. GTNs do not pose any limitations on the model architecture. NFs also differ from GTNs in their optimization method since NFs employ a maximum likelihood objective – typically the Kullback-Leibler (KL) divergence while GTNs are trained using MSE. It was observed that optimizing the KL-divergence may be difficult for more complex normalizing flow structures (Xu, Chen, and Campbell 2023). NFs may also suffer from low-quality sample generation (Behrmann et al. 2021), partially because the constraint on the Jacobian, besides limiting the model structure, may also lead to issues such as spurious local minima and numerical instability (Golinski, Lezcano-Casado, and Rainforth 2019).
@@
Continuous normalizing flow (CNF) are a relatively recent type of NFs that were developed to avoid the main constraints posed by NFs – namely the requirements of invertibility and a tractable Jacobian determinant. To avoid these constraints, CNFs use ordinary differential equations (ODEs) to describe the transformations in NFs. While CNFs have provided certain improvements over NFs, they are still time-intensive since they require solving a series of ODEs, equivalent to running a neural network with hundreds of layers (Huang and Yeh 2021). Improving both the speed and performance of CNFs is an active and promising field of research. A comparison between a recent CNF and GTN was suggested in the Experiments section.
@@ 5 Discussion
In this work we introduced a new class of generative models – GTNs – that learn a homeomorphism $h$ between the supports of two continuous random variables using samples from both. The samples are used in an efficient algorithm to create a training dataset for the GTN which is trained in a simple supervised approach, with no restrictions on the model architecture. The freedom to use a standard neural network architecture, and the fact that $h$ satisfies the conditions of the universal approximation theorem, lead GTNs to quickly converge towards $h$, providing an efficient method to generate samples from a target random variable. This was demonstrated experimentally in multiple cases, particularly manifesting in the observation that generated samples reached the data manifold at early epochs and in the total time to convergence. Other positive implications of $h$ being a homeomorphism – namely, coverage and diversity, continuous interpolations and guiding topological properties – were also discussed and demonstrated experimentally.
One downside of our approach is that it relies on the theoretical assumption that the data is represented in its intrinsic dimension. Experimentally, this assumption manifested in different performances and sample quality across different dimensions. This means that dimensionality reduction methods, like the autoencoders used in this work, may be essential for high quality sample generation. Future work, e.g. for generating high-resolution images, may be interested in experimenting with stronger autoencoders as well as additional procedures such as data augmentations, which were not used here. Additionally, the assumptions on the supports of the distributions mean that our method may be ill-suited for discrete distributions, and that datasets may need to be split into separate components for individual training, as explained in Appendix C. The latter case means that larger datasets may be required to ensure that the number of samples per component is sufficiently large. Existing literature on estimating the intrinsic dimension of the data and on the relation between the intrinsic dimension and the required sample complexity (e.g., (Pope et al. 2021)) may be of value in both of these contexts.
@@
To understand the rationale behind Algorithm 1, imagine four points from each of $X$ and $Y$ lying on the same line, say the $x$-axis in $\mathbb{R}^{2}$, such that both $X$ and $Y$ have two points on each side of the origin (we assume $D_{X},D_{Y}$ are centralized). See Appendix Figure 7, where circles are from $Y$ and stars are from $X$. The cosine similarity of each pair of points $\{a,b\}$ with $a\in D_{X},b\in D_{Y}$ is $1$ if they are on the same side or $-1$ of they are on opposite sides. Take the first $y\in D_{Y}^{\text{sorted}}$ (the closest one to the origin which reflects $\sim 50$th percentile since we assumed an equal number of points from $Y$ on both sides). The maximum cosine similarity out of all $x\in D_{X}^{\text{sorted}}$ is $1$, meaning $y$ will be matched with an $x$ on the same side. The fact that $D_{X}$ is sorted, means that we are using the distance as a tie breaker – out of all $x\in D_{X}$ that are on the same side as $y$, the $x$ that is closest to the origin ($\sim 50$th percentile) will be chosen as $x_{y}$. Likewise, the leftmost $y$ will be matched with the leftmost $x$ (both 00th percentile), the rightmost $y$ with the rightmost $x$ (both $100$th percentile) etc. Note that if this were in $\mathbb{R}$, this process coincides with the same labeling process described in the $1$D case (illustrated in Figure 2).
@@
One example of a topological property is the connectedness property – being a ”single piece”, like $\mathbb{R}$, as opposed to ”more than one piece”, like $(-\infty,-1)\cup(1,\infty)$, is preserved between homeomorphic spaces (so, in particular, these two are not homeomorphic, but $\mathbb{R}$ is homeomorphic to each of the two intervals, separately). Therefore, if we suspect that our data is composed of separate classes like images of both hands and faces, and we would like to train a GTN as a generative model, we could split our data to its different components (learning a separate $h$ for each). This relates to the conditions of Theorem 2.1 which assume that $X$’s support is an open interval. For example, if we suspect that our data can be separated into two disjoint uniform distributions, then the theorem doesn’t apply. However, it does apply to each separate uniformly distributed component.
@@
Figure 9: Inception score (IS) vs real IS for CelebA plotted by epoch (at epochs of improvement in the IS). At each epoch, the IS was computed for $200$ randomly generated images (decoded latent vectors generated by $\hat{h}$ from random normal vectors, formally: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ for $r\sim\mathcal{N}(\textbf{0},\textbf{I})$) (solid black line). Real IS was computed once for a random set of $200$ real images (decoded latent vectors of real images) (dashed gray line).
Figure 10: Generated vs real samples for latent dimension $d=10$. (A) Random generated samples – each image is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). (B) Random real decoded images – each image is the decoded latent vector of a random real image.
Figure 11: Generated vs real samples for latent dimension $d=25$. (A) Random generated samples – each image is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). (B) Random real decoded images – each image is the decoded latent vector of a random real image.
Figure 12: Generated vs real samples for latent dimension $d=50$. (A) Random generated samples – each image is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). (B) Random real decoded images – each image is the decoded latent vector of a random real image.
Figure 13: Generated vs real samples for latent dimension $d=100$. (A) Random generated samples – each image is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). (B) Random real decoded images – each image is the decoded latent vector of a random real image.
@@
Figure 16: Inception score (IS) vs real IS for HaP plotted by epoch (at epochs of improvement in IS). At each epoch, IS was computed for $200$ randomly generated images (decoded latent vectors generated by $\hat{h}$ from random normal vectors), formally: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ for $r\sim\mathcal{N}(\textbf{0},\textbf{I})$) (solid black line). Real IS was computed once for a random set of $200$ real images (decoded latent vectors of real images) (dashed gray line).
@@
Final architecture specifications for $\hat{h}$ appear in Appendix Table 1. We ran a hyperparameter search for CelebA and HaP. Initially, we ran several settings for CelebA using a small number of epochs (roughly 100-300) for latent dimension $200$ prior to consulting the literature on the intrinsic dimension of image data. We tried learning rates of $1e^{-3},1e^{-4},1e^{-5},5e^{-5}$ for various width and depth settings. Specifically, for CelebA we tested widths of $500,1000,1200,1300,1500$ and depths of $10,17,20,25,27$. For HaP we kept the depth at $25$ after observing this obtained best results for CelebA and tested depths of $2000,3000$ as well after identifying that HaP benefited from higher width settings.
@@
Table 1: Architecture specifications for $\hat{h}$ for the different datasets. The architecture was kept the same between the different dimensions where relevant (CelebA, HaP).
@@
IS $\uparrow$ IS real FID $\downarrow$ IS $\uparrow$ IS real FID $\downarrow$
$d=10$ 1.97 1.95 36.66 1.90 1.80 81.50
$d=25$ 1.93 1.99 32.11 2.63 2.43 88.85
$d=50$ 1.88 2.06 33.84 2.64 2.22 88.34
$d=100$ 1.86 2.07 36.43 2.37 1.94 124.62
Table 2: Results per dataset and latent dimension. Inception Score (IS) is the best IS attained for the given dimension and dataset for a random set of 200 generated images. IS real is the IS attained by a set of 200 random real images (decoded latent vectors of real encoded images). Fréchet Inception Distance (FID) is computed for the same model that attained the reported IS using all training images and a same-sized set of randomly generated images.

### v3

@@ 4 Related work
GTNs are related in concept to NFs which seek to map a source distribution into a target distribution using a sequence of bijective transformations. These are typically implemented by a neural network, often in the form of at least tens of neural network blocks (Xu and Campbell 2024) and sometimes more (Xu, Chen, and Campbell 2023). The core of NFs is based on each of these transformations being a diffeomorphism – a specific type of homeomorphism that is more constrained than $h$ as it is defined between smooth manifolds (as opposed to any topological spaces, including any manifolds) and requires that the function and its inverse are differentiable (as opposed to just continuous for general homeomorphisms). NFs also require that the log-determinant of the Jacobian of these transformations is tractable, posing a constraint on the model architecture. GTNs do not pose any limitations on the model architecture. NFs also differ from GTNs in their optimization method since NFs employ a maximum likelihood objective – typically the Kullback-Leibler (KL) divergence while GTNs are trained using MSE. It was observed that optimizing the KL-divergence may be difficult for more complex normalizing flow structures (Xu, Chen, and Campbell 2023). NFs may also suffer from low-quality sample generation (Behrmann et al. 2021), partially because the constraint on the Jacobian, besides limiting the model structure, may also lead to issues such as spurious local minima and numerical instability (Golinski, Lezcano-Casado, and Rainforth 2019).
@@
Continuous normalizing flow (CNF) are a relatively recent type of NFs that were developed to avoid the main constraints posed by NFs – namely the requirements of invertibility and a tractable Jacobian determinant. To avoid these constraints, CNFs use ordinary differential equations (ODEs) to describe the transformations in NFs. While CNFs have provided certain improvements over NFs, they are still time-intensive (Huang and Yeh 2021). Improving both the speed and performance of CNFs is an active and promising field of research.
Also closely related are VAEs since they are based on autoencoders that aim to transform the lower-dimensional latent space into a tractable distribution. The VAE loss function optimizes two terms, namely both the reconstruction error and the error between the prior and posterior distributions. This often leads to training instabilities and to posterior collapse. Despite these training instabilities, methods based on VAEs have been suggested, including Two-Stage VAE (Dai and Wipf 2019) which first trains a VAE to identify the lower-dimensional manifold, and then a second VAE to transform the learned latent space into a normal distribution. Other methods, combining VAE and flow, have also been suggested, but have seen limited success at improving upon VAE (Xiao, Yan, and Amit 2019) and add architectural limitations as mentioned earlier. Compared to VAE-based methods, GTNs have several practical advantages. One advantage is that GTNs separate between the autoencoder and the generative process, avoiding the intricacies of balancing between the two terms present in the VAE loss function (an in-depth analysis of the issues arising from this balancing can be found in (Dai and Wipf 2019)). Practically, this means that one can focus on finding a high-quality latent representation of the data first, and then learn to generate in that latent space. This may be of particular interest in light of the use of pretrained autoencoders in state-of-the-art generative models (Rombach et al. 2022). Another advantage is the training stability, with no risk of posterior collapse, which is a main issues with VAEs. Also worth noting is the fact that GTNs show no evidence of mode collapse (a known issue with other methods like GANs). These advantages, and the fact that GTNs employ a standard supervised learning approach, makes them more user-friendly and easier to train than many of the existing methods.
@@ 5 Discussion
This work provides two main contributions: (1) It introduces a new class of generative models – GTNs – that offers training stability and simplicity; (2) It provides a new perspective on the importance of the intrinsic dimension of the data in the context of generative models.
As a generative method, GTNs offer a simple and stable approach to image generation. Compared to other generative methods like diffusion, VAE, GANs and NFs, GTNs are computationally simpler and more user-friendly since they employ a basic supervised learning approach that does not suffer from many of the training instabilities and specialised architectural requirements posed by other methods. GTNs also allow the flexibility of easily replacing the autoencoder should a better one be designed. This is of value in light of the recent turn to using pre-trained autoencoders for data generation (Rombach et al. 2022). VAEs, for example, would require retraining from scratch, which means risking dealing again with training instabilities and potential posterior collapse.
GTN’s potential was demonstrated both quantitatively and qualitatively, particularly with demonstrable improvements over VAE. It is possible that with careful fine-tuning and with more sophisticated architectures to replace the fully-connected GTNs (e.g. to convolutional-based GTNs), further improvements may be attained (especially for CIFAR-10 where no attempts were made to fine-tune the settings or adapt the architecture when transitioning from CelebA). Upgrading from the vanilla autoencoders may also offer improvements, perhaps even more significant ones – observing both the qualitative and quantitative similarity between reconstructions and generated samples (e.g. recon-IS being nearly achieved, the similarity between FID and real-FID as well as the similarities observed in Figure 6) shows that a large portion of the error may be attributed to the quality of the autoencoder.
GTNs can also be extended to data that does not immediately satisfy the assumptions on the support of the distribution (like data that is separable into disconnected components). In such a case, it is possible to exploit the assumption that the data lies on a manifold by labeling points ’locally’ using a mixture model before training the GTN. We demonstrate this in Figure 9 where the data distribution is supported on two disjoint sets. Note that GTNs were also successfully applied to distributions that are not rotation invariant, particularly the uniform distribution. We leave a more thorough discussion and analysis of such extensions to future work.
Figure 9: Generating two disjoint uniform distributions. Left-to-right: mixture of random normal samples, corresponding predictions generated by a GTN and actual samples from the data. The random normal samples were generated by clustering the training data into two clusters using the fast-pytorch-kmeans package (DeMoriarty 2024) and using the cluster means and standard-deviations to define the normal distributions. Labels were computed in each cluster separately according to Algorithm 1 to obtain a single dataset for training the GTN.
Finally, the topological considerations used throughout this paper, including those relating to the intrinsic dimension of the data, may prove useful for other methods, as already briefly discussed for diffusion. For example, NFs are by-definition sensitive to the same topological considerations discussed herein since they employ diffeomorphisms – a strict form of homeomorphisms. In particular, this means that topological consideration like connectedness (Appendix C) and the intrinsic dimension are important to consider there too. We note that particularly the intrinsic dimension of the data is an important property that is largely overlooked – many generative methods have been used in the higher-dimensional pixel-space. Recently, latent generative methods have been developed and were shown to improve upon higher-dimensional ones (Rombach et al. 2022). One potential contribution to that success could lie in the fact that the latent dimension is closer to the intrinsic dimension of the data, reducing the risk of generating many outliers like those produced by diffusion in the provided swiss-roll example. More broadly, topological considerations like those used throughout this paper may prove useful for the design of better generative models and for the understanding of why some methods improve upon others.
@@
To understand the rationale behind Algorithm 1, imagine four points from each of $X$ and $Y$ lying on the same line, say the $x$-axis in $\mathbb{R}^{2}$, such that both $X$ and $Y$ have two points on each side of the origin (we assume $D_{X},D_{Y}$ are centralized). See Appendix Figure 10, where circles are from $Y$ and stars are from $X$. The cosine similarity of each pair of points $\{a,b\}$ with $a\in D_{X},b\in D_{Y}$ is $1$ if they are on the same side or $-1$ of they are on opposite sides. Take the first $y\in D_{Y}^{\text{sorted}}$ (the closest one to the origin which reflects $\sim 50$th percentile since we assumed an equal number of points from $Y$ on both sides). The maximum cosine similarity out of all $x\in D_{X}^{\text{sorted}}$ is $1$, meaning $y$ will be matched with an $x$ on the same side. The fact that $D_{X}$ is sorted, means that we are using the distance as a tie breaker – out of all $x\in D_{X}$ that are on the same side as $y$, the $x$ that is closest to the origin ($\sim 50$th percentile) will be chosen as $x_{y}$. Likewise, the leftmost $y$ will be matched with the leftmost $x$ (both 00th percentile), the rightmost $y$ with the rightmost $x$ (both $100$th percentile) etc. Note that if this were in $\mathbb{R}$, this process coincides with the same labeling process described in the $1$D case (illustrated in Figure 2). Note that if $D_{X},D_{Y}$ are not sorted, discontinuities in the labeling may arise (see Figure 11).
@@
Figure 11: Example demonstrating that sorting $D_{X},D_{Y}$ in the labeling algorithm is essential. Training GTN on the 2D-uniform data on labels produced by the same algorithm with sorting removed causes test predictions to collapse mainly to the boundary.
@@
One example of a topological property is the connectedness property – being a ”single piece”, like $\mathbb{R}$, as opposed to ”more than one piece”, like $(-\infty,-1)\cup(1,\infty)$, is preserved between homeomorphic spaces (so, in particular, these two are not homeomorphic, but $\mathbb{R}$ is homeomorphic to each of the two intervals, separately). Therefore, if we suspect that our data is composed of separate classes like images of both hands and faces, and we would like to train a GTN as a generative model, one option would be to use a clustering approach and to generate labels for each cluster separately. Another would be to split our data into its different components (learning a separate $h$ for each). This relates to the conditions of Theorem 2.1 which assume that $X$’s support is an open interval. For example, if we suspect that our data can be separated into two disjoint uniform distributions, then the theorem doesn’t apply. However, it does apply to each separate uniformly distributed component.
@@
Figure 13: Inception score (IS) vs recon-IS for CelebA plotted by epoch (at epochs of improvement in the IS). At each epoch, the IS was computed for $200$ randomly generated images (decoded latent vectors generated by $\hat{h}$ from random normal vectors, formally: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ for $r\sim\mathcal{N}(\textbf{0},\textbf{I})$) (solid black line). recon-IS was computed once for a random set of $200$ reconstructions (dashed gray line).
Figure 14: Inception score (IS) vs recon-IS for HaP plotted by epoch (at epochs of improvement in IS). At each epoch, IS was computed for $200$ randomly generated images (decoded latent vectors generated by $\hat{h}$ from random normal vectors), formally: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ for $r\sim\mathcal{N}(\textbf{0},\textbf{I})$) (solid black line). recon-IS was computed once for a random set of $200$ reconstructions (dashed gray line).
@@
Final architecture specifications for $\hat{h}$ appear in Appendix Table 2. We ran a hyperparameter search for CelebA and HaP. Initially, we ran several settings for CelebA using a small number of epochs (roughly 100-300) for latent dimension $200$ prior to consulting the literature on the intrinsic dimension of image data. We tried learning rates of $1e^{-3},1e^{-4},1e^{-5},5e^{-5}$ for various width and depth settings. Specifically, for CelebA we tested widths of $500,1000,1200,1300,1500$ and depths of $10,17,20,25,27$. For HaP we kept the depth at $25$ after observing this obtained best results for CelebA and tested widths of $2000,3000$ as well after identifying that HaP benefited from higher width settings.
@@
Appendix G
The 1D swiss-roll data is generated by sampling $\theta\sim U(1.5\pi,4.5\pi)$ and computing $f(\theta)=\theta(\cos{\theta},\sin{\theta})$. GTN is trained to generate $\theta$. This example, besides demonstrating the 1D case, serves to demonstrate how GTN operates on the ’correct’ latent representation of the data.
Appendix H
@@
CIFAR-10 26 1,500
LeakyReLU(0.5)
Yes $5e^{-5}$ No Adam 200
Table 2: Architecture specifications for $\hat{h}$ for the different datasets. The architecture was kept the same between the different dimensions where different dimensions were tested (CelebA, HaP).
@@
IS $\uparrow$ recon-IS $\uparrow$ FID $\downarrow$ IS $\uparrow$ recon-IS $\uparrow$ FID $\downarrow$
$d=10$ 1.97 1.95 119.46 1.90 1.80 192.98
$d=25$ 1.93 1.99 82.88 2.63 2.43 123.89
$d=50$ 1.88 2.06 71.61 2.64 2.22 81.48
$d=100$ 1.86 2.07 66.05 2.37 1.94 101.85
Table 3: Results per dataset and latent dimension using vanilla autoencoders (2 layer encoder). Inception Score (IS) is the best IS attained during training for the given dimension and dataset for a random set of 200 reconstructions. recon-IS is the IS attained by a set of 200 reconstructions. Fréchet Inception Distance (FID) is computed using $50,000$ randomly generated images and $50,000$ random real images using the same model that attained the reported best IS. For the smaller HaP dataset, all training images ($\sim$ 8,000) and the same number of randomly generated images were used.
Method FID $\downarrow$
2-Stage VAE (Dai and Wipf 2019) 44.4
NVAE (Vahdat and Kautz 2020) 14.74
WAE-GAN (Tolstikhin et al. 2017) 42.0
DiffuseVAE (Pandey et al. 2022) 3.97
GLF (Xiao, Yan, and Amit 2019) 53.2
VAE + flow posterior (Xiao, Yan, and Amit 2019) 67.9
VAE + GMM=75 (Pandey et al. 2022) 72.11
GTN (ours) 66.05
Table 4: GTN results and results reported by related methods on CelebA. GTN results are for $d=100$ (from Table 3). All other results are as originally reported by their authors, except for NVAE which does not report FID (results are as reported by (Pandey et al. 2022)). Note that the methods may differ in their evaluation settings and implementation.
Method FID $\downarrow$
2-Stage VAE (Dai and Wipf 2019) 72.9
NVAE (Vahdat and Kautz 2020) 51.67
DiffuseVAE (Pandey et al. 2022) 2.62
GLF (Xiao, Yan, and Amit 2019) 88.3
VAE + flow posterior (Xiao, Yan, and Amit 2019) 143.6
VAE + GMM=75 (Pandey et al. 2022) 137.68
GTN (ours) 238.62
Table 5: GTN results and results reported by related methods on CIFAR-10. GTN results are from Table 1. All other results are as originally reported by their authors, except for NVAE which does not report FID (results are as reported by (Pandey et al. 2022)). Note that the methods may differ in their evaluation settings and implementation.
Figure 17: Generated images by: (left) decoding random normal vectors with vectors $\mu,\sigma$ estimated from the latent space (each image is the decoded $r$ for some random $r\sim\mathcal{N}(\mu,\sigma I)$); (right) GTN (each image is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(0,I)$)).